How are returns on common stocks in overseas markets related to returns in one countries markets? Consider measuring the countries returns by the annual rate of return on the index A and overseas returns by the annual rate of return on index B. Both are recorded in percents. We will regress the B returns on the A returns for 21 years. Here is part of the output for this regression. The regression equation is

B = −3.17 + 0.811A.

Complete the analysis of variance table by filling in the missing boxes. (Round your answer for F to two decimal places and your answers for SS and MS to one decimal place.)

Analysis of Variance
Source DF SS MS F
Regression 1 5552.1
Residual Error
Total 10077.1
Answer #1

TOPIC:Finding the missing values of the ANOVA TABLE.

pegreselon model-2.1 i-e, po andß k-1 = 2-1 =1 So, Regnus ion dt CEonor) 二1 201. Regression 5552 5552:1 Residual (Ronon)ss Total ss-Regxuld :on 1007 -55521 4525 Residual ss Resi dual Re亼 4192 5EL23의Regressiom Ms Residual Ecor ms 0 5552 238 2 二 123-31 So, the requina ed ANovA TABLE s→ m S 5552 5552 23.31 Regrexsion Residual (Enno 1g 4525 238.2 20 O07.I -Totaレ Ans